Can someone help me solve this? I couldnt find it in my textbook and I am confused.

Thanks!Cherry Blossom Products Inc. produces and sells yoga-training products: how-to DVDs and a basic equipment set (blocks, strap,1. What is the sales mix of DVDs and equipment sets? Select 2. Compute the break-even quantity of each product. The break-eve
Answer #1

DVDs = 8,960

Equipment sets = 4,480

Total = 8,960 + 4,480 = 13,440

1. Sales mix

DVDs = 8,960 / 13,440

= 66.67% or 2/3

Equipment sets = 4,480 / 13,440

= 33.33% or 1/3

2. Contribution margin per unit = Selling price per unit - Variable costs per unit

DVDs = $8.1 - $4.5 = $3.6

Equipment sets = $25.2 - $14.9 = $10.3

Weighted average contribution margin per unit = (DVDs contribution margin per unit * Sales mix) + (Equipment sets contribution margin per unit * Sales mix)

= ($3.6 * 2/3) + ($10.3 * 1/3)

= $2.4 + $3.43

= $5.83

Break-even quantity = Fixed costs / Weighted average contribution margin per unit

= $66,500 / $5.83

= 11,407 units

Break-even quantity for DVDs = 11,407 * 2/3 = 7,605

Break-even quantity for equipment sets = 11,407 * 1/3 = 3,802

* Break-even quantity is rounded to 0 decimal places.